OpenStudy (anonymous):

Shruthi moves along a circular track of radius 7 m starting from point A.Find the displacement and distance when she reaches the point a)At B b)At C c)At D d)At A again? A B C D Reply Using Drawing

OpenStudy (anonymous):



OpenStudy (abhisar):

Displacement: 

OpenStudy (abhisar):

When shruthi reach B, displacement will be equal to x. By pythagorus theorem \(\large \sf x=\sqrt{7^2+7^2}=\sqrt{2}\times7m\)

OpenStudy (abhisar):

When shruthi reaches point C, then displacement=2.R=2*7=14m. Again when she reaches D, displacement=y which is equal to x =\(\large \sf 7\sqrt{2}\)m

OpenStudy (abhisar):

When shruthi reaches back to A, then displacement = 0 !

OpenStudy (abhisar):

Distance: At point A => \(\large \sf \frac{2\pi\times7}{4}=11m\)

OpenStudy (abhisar):

Distance: At point B => \(\large \sf \frac{2\pi\times7}{4}=11m\) At point C => 2*7=14m At point D => \(\sf \large \frac{3\times2\pi\times7}{4}=?\) Back to point A => \(\large \sf 2\pi \times 7=44m\)

OpenStudy (anonymous):

@Abhisar At point C both distance and displacement is 14 m ?

OpenStudy (anonymous):

@aryandecoolest help here ..

OpenStudy (anonymous):

no its not @aryandecoolest

OpenStudy (anonymous):

its confusing

OpenStudy (anonymous):

see if you are confused about point C then since from A to C it covers semi circle, so distance covered at C will be =\[22/7 *r=22/7*7=22m\] and displacement will be 7+7=14m

OpenStudy (anonymous):

distance will be equal to the displacement and the speed equal to the velocity when the motion is along the straight line, and it remains so also the magnitude has to remain constant which means that it'll be travelling with uniform velocity. ... Now since direction changes , so it's not possible for object in circle to have Distance=Displacement...

OpenStudy (abhisar):

Yes, there will be a rectification \(\color{blue}{\text{Originally Posted by}}\) @Abhisar Distance: At point B => \(\large \sf \frac{2\pi\times7}{4}=11m\) At point C => pi*7=22m At point D => \(\sf \large \frac{3\times2\pi\times7}{4}=?\) Back to point A => \(\large \sf 2\pi \times 7=44m\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

is it clear now ?

OpenStudy (anonymous):

how 7 +7 ?

OpenStudy (anonymous):

@Abhisar and @aryandecoolest how 7+7 ?

OpenStudy (abhisar):

You mean how displacement is 14 at c ?

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

Displacement between two points = shortest distance between that two points Shortest distance between A and C will be equal to the diameter of the circle. Since radius is 7, displacement = 7+7 [or 7*2]=14

OpenStudy (anonymous):

It is a straight line connecting A and C....which is the shortest path. so it's a diameter which connects A and point C, opposite to A.

OpenStudy (anonymous):

ok thats clear now . Sorry i last doubt how the displacement will be 0 of point D to A ?

OpenStudy (abhisar):

Displacement will be 0 at point a and will be 7\(\sqrt{2}\)

OpenStudy (abhisar):

at point D

OpenStudy (anonymous):

sorry, I didn't get u :(

OpenStudy (abhisar):

When u come back to ur initial position then ur displacement is counted as zero. Shruthi started at point A so her displacement will be counted as 0 if she comes back to point A

OpenStudy (abhisar):

At point D, the shortest distance between A and D will be equal to the arc AD and hence displacement will be 7\(\sqrt{2}\)

OpenStudy (anonymous):

ok now i get it thank u so much and sorry for disturbing

OpenStudy (abhisar):

It's allryt :) Happy to help

OpenStudy (anonymous):

@aryandecoolest to you too thanks for helping me :)

OpenStudy (anonymous):

anytime)

OpenStudy (anonymous):

@aryandecoolest sorry for asking this foolish question are u a boy or girl?

OpenStudy (anonymous):

LOL @Greaty what you think i may be? :)

OpenStudy (anonymous):

of course I know that u r a human...:)

OpenStudy (anonymous):

hahaha wow you are genius @Greaty OF course i am human not alien but in that what's my gender??

OpenStudy (anonymous):

A GIRL?

OpenStudy (anonymous):

LOL sorry you went complete wrong @Greaty

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

never mind.... :) :P

OpenStudy (anonymous):

anyway u a great friend. U belongs to which place?